### 14 Mar 2026 v7

@@ 1 Introduction
In Chapter 3, we propose a novel encoding mechanism which we named Eagle. For the same key parameters, the Eagle encoding algorithm encodes the same input text sequence into completely different output text sequences by introducing randomness, and the output sequence can be considered completely random. The distinguishing feature of Eagle encoding algorithm from any other encoding algorithm is that for any given group of input and output texts, the key parameters are completely unsolvable, meaning that any key parameter can solve for the same input and output text.
@@
In fact the encryption algorithm developed in Chapter 4 does not meet the security requirements for the key. This is because, when knowing the plaintext-ciphertext correspondence, although the complexity of cracking the key is exponential to the length of the key, it does not guarantee that only exhaustive search is effective for cracking the key. There may be some other attack methods that can effectively reduce the difficulty of cracking the key. Such as for a 256 bit key, although it can ensure that the computational complexity of cracking the key is at least $O(2^{64})$, this is still far from the ideal security of $O(2^{256})$. The encryption algorithm constructed in Chapter 4 only analyzes the computational complexity of cracking the key given a plaintext ciphertext correspondence, which does not guarantee its ability to resist linear and differential attacks, For attackers, they can construct plaintexts with certain characteristics multiple times, then find some structures that we did not anticipate from the relationship between plaintext and ciphertext.
In Chapter 5, we first delve into the mathematical laws behind the Eagle encoding algorithm and discover an operation that satisfies the allocation law for XOR operations. This operation is very similar to the allocation law of multiplication operations, but does not satisfy the commutative and associative laws, we refer to it as a “right multiplication operation on binary”. If we consider this right multiplication operation as a linear operation of one parameter on another, we can easily construct nonlinear operations of one parameter on another, such as “high-order” operations. Based on this nonlinear construction, we redesigned a block symmetric encryption algorithm and named it $Eagle^{*}$. For $Eagle^{*}$, knowing any plaintext ciphertext correspondence, there is no more effective way for attackers to crack the key than exhaustive search.
@@ 2 Notations
$\mathbb{E}_{w_{0},w_{1}}(s_{0},m)$ - Excute Eagle encoding algorithm as shown in Algorithm 1.
$\mathbb{D}_{w_{0},w_{1}}(c,s_{L})$ - Excute Eagle decoding algorithm as shown in Algorithm 2.
$m*\delta$ - Excute the right multiplication of $m$ with $\delta$ introduced in chapter 5.2.
$m*\delta^{2}$ - Excute the right multiplication of $m$ with $\delta$ twice as $m*\delta*\delta$.
$s/\delta$ - Excute the right division of $s$ to $\delta$ introduced in chapter 5.2.
$s/\delta^{2}$ - Excute the right division of $s$ to $\delta$ twice as $s/\delta/\delta$.
@@ 3 Introduction to Eagle encoding algorithm
$\mathbb{E}_{w_{0},w_{1}}(m,s_{0})=(c,s_{L})$ : use the key $(w_{0},w_{1})$ to execute Eagle encoding algorithm (Algorithm 1) on any initial state $s_{0}$ and input $m$ to obtain $c$ and $s_{L}$.
$\mathbb{D}_{w_{0},w_{1}}(c,s_{L})=(m,s_{0})$ : use the key $(w_{0},w_{1})$ to execute Eagle decoding algorithm (Algorithm 2) on any $c$ and $s_{L}$ to obtain $s_{0}$ and $m$.
@@
Next, we introduce another useful conclusion about the Eagle encoding algorithm.
[Theorem 3] For any equation $\mathbb{E}_{w_{0},w_{1}}(m,*)=(c,*)$, which $(w_{0},w_{1})$ are unknown variables and $m,c$ are known variables, this equation can be seen as an identity, valid for all $(w_{0},w_{1})$.
Proof.
This theorem requires us to construct an algorithm that, given any $w_{0},w_{1}$ to find a valid equation $\mathbb{E}_{w_{0},w_{1}}(m,*)=(c,*)$.
When given $w_{0},w_{1}$, randomly choose $s_{0}^{{}^{\prime}}$, excute $\mathbb{E}_{w_{0},w_{1}}(m,s_{0}^{{}^{\prime}})=(c^{{}^{\prime}},s_{L})$, according to the theorem 2, $s_{L}$ is irrelevant to $s_{0}^{{}^{\prime}}$, then we excute $\mathbb{D}_{w_{0},w_{1}}(c,s_{L})=(m,s_{0})$, thus $\mathbb{E}_{w_{0},w_{1}}(m,s_{0})=(c,s_{L})$. That is to say, when only $m,c$ is known, any $w_{0},w_{1}$ is valid.
If we construct an encryption algorithm where $(w_{0},w_{1})$ is the key, $m$ is the plaintext, $c$ is the ciphertext, $s_{0}$ and $s_{L}$ never appear in the plaintext sequences or the ciphertext sequences, according to the conclusion of Theorem 3, for a cracker, under the condition of only knowing the plaintext ciphertext correspondence $m,c$, any value $(w_{0},w_{1})$ in the key space is valid and is indistinguishable, this feature allows us to construct the security that meets the requirements of computational complexity.
4 Construction of one-way functions
4.1 Introduction to one-way functions
[Theorem 4] If one-way function exists, then $P\neq NP$.
4.2 Construction of one-way functions
In this chapter, we constructed a short key symmetric encryption algorithm, which we named it Eagle. The key consists of four independent variables, each of which contains at least $2^{L-1}$ possible values. During the encryption process, several intermediate random numbers are generated at the encryption party, after the encryption process is completed, these intermediate random numbers are directly deleted at the encryption party, and only the ciphertext is sent to the decryption party. Under the condition of knowing the correspondence between any multiple sets of plaintext-ciphertext, the problem of cracking the key is equivalent to solving a system of equations. In the process of solving this system of equations, at least one free unkown variables cannot be eliminated, so it is necessary to traverse all possible values of at least this variable, which means that the computational complexity for cracking the key is at least $O(2^{L})$.
It should be noted that although the computational complexity of cracking the key is exponentially under the condition of knowing the correspondence between plaintext and ciphertext, this does not mean that this encryption algorithm is safe and practical. Assume the length of the key is 4L24L-2 bits, although the computational complexity of cracking the key is not less than $O(2^{L})$, may be much lower than the expected complexity of $O(2^{4L-2})$. A secure and practical encryption algorithm needs to ensure that the cracking key can only exhaustively search for every possible key which is the issue to be considered in the next chapter. In this chapter, our goal is to design a one-way function where we only need to design and prove that the computational complexity for cracking keys is exponential to the length of the key.
4.2.1 $Eagle$ key generator
The key is $(w_{0},w_{1})$ and $(w_{2},w_{3})$, where $w_{0}$ and $w_{1}$ have odd number of bits different, $w_{2}$ and $w_{3}$ have odd number of bits different.
Assuming $w_{0},w_{1},w_{2},w_{3}$ are all $L$ bits, the length of the entire key can be considered as 4L24L-2 bits, this is because they have a loss of 2 bit.
We randomly generate a $2L$ bit number, taking the first $L$ bit as $w_{0}$. When the last $L$ bit has an odd number of bits that are different from $w_{0}$, take the last $L$ bit data as $w_{1}$. When the last $L$ bit has an even number of bits that are different from $w_{0}$, take the last bit inverted to the last $L$ bit as $w_{1}$.
For example, we generate a number with 16 bits as $1001001101001000$, we set $w_{0}=10010011$ (the first 8 bits) and $w_{1}=01001000$ (the last 8 bits), when $w_{0}$ and $w_{1}$ have 6 (even) bits different, we set the last bit of $w_{1}$ inverted as $w_{1}=01001001$.
For $(w_{0},w_{1})$, because the probability of odd bits different are the same as even bits different, the total number of $(w_{0},w_{1})$ is $2^{2L-1}$.
For $(w_{2},w_{3})$, the generation method is the same as $(w_{0},w_{1})$, the total number of $(w_{2},w_{3})$ is also $2^{2L-1}$.
4.2.2 $Eagle$ encryption process
The plaintext $P$ is grouped by L1L-1 bits, and the last group with less than L1L-1 bits are randomly filled into L1L-1 bits. The total number of groups is assumed to be $K$, denoted as $P=(P_{1},P_{2},...,P_{K})$.
Then, one bit is randomly filled into the last bit of each $P_{i}(1\leq i\leq K)$, making the length of $P_{i}$ equal to $L$.
Generate random numbers of $K$ groups with $L$ bits completely independently, denoted as $(x_{1},x_{2},...,x_{K})$.
Generate another three random numbers denoted as $S_{0}$, $S_{0}^{{}^{\prime}}$ and $y_{1}$.
The encryption process is described in the following algorithm.
Algorithm 3 EagleEagle-Encrypt.
0:Key: $(w_{0},w_{1})$, $(w_{2},w_{3})$. Plaintext sequence: $P=(P_{1},P_{2},...,P_{K})$. Random number sequence: $(x_{1},x_{2},...,x_{K})$. Random number: $S_{0},S_{0}^{{}^{\prime}},y_{1}$.
0:Ciphertext sequence: $(C_{1},C_{2},...,C_{2K})$ and $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$.
[…]
2:for $i=1;i\leq K;i++$ do
3:  $(C_{2i-1},S)\leftarrow\mathbb{E}_{w_{0},w_{1}}(P_{i},S)$
4:  $S\leftarrow(S\oplus(1\&x_{i}))$ {The last bit of $S$ needs to be XORed with the last bit of $x_{i}$ to hide the last bit of $S$ from $C_{2i}$}
5:  $(C_{2i},S)\leftarrow\mathbb{E}_{w_{0},w_{1}}(x_{i},S)$
6:  if $i<K$ then
7:   $S\leftarrow(S\oplus(1\&P_{i+1}))$ {The last bit of $S$ needs to be XORed with the last bit of $P_{i+1}$ to hide the last bit of $S$ from $C_{2i+1}$}
8:  end if
9:end for
[…]
The above algorithm can be represented by the following figure.
$S_{0}$$C_{1},S_{1}$$C_{2},S_{2}$$w_{0},w_{1}$$C_{2},S_{2}$$C_{3},S_{3}$$C_{4},S_{4}$$C_{4},S_{4}$$C_{5},S_{5}$$C_{6},S_{6}$$w_{2},w_{3}$$S_{0}^{{}^{\prime}}$$C_{1}^{{}^{\prime}},S_{1}^{{}^{\prime}}$$C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}}$$\scriptstyle{P_{1}}$$\scriptstyle{x_{1}}$$\scriptstyle{P_{2}}$$\scriptstyle{x_{2}}$$\scriptstyle{P_{3}}$$\scriptstyle{x_{3}}$$\scriptstyle{y_{0}=S_{6}}$$\scriptstyle{y_{1}}$
4.2.3 $Eagle$ decryption process
For the decrypting party, the key is also $(w_{0},w_{1}),(w_{2},w_{3})$, the ciphertext is $(C_{1}$, $C_{2}$, …, $C_{2K})$ and $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$. Please refer to the following for details.
Algorithm 4 EagleEagle-Decrypt.
0:Key: $(w_{0},w_{1}),(w_{2},w_{3})$. Ciphertext sequence: $(C_{1},C_{2},...,C_{2K})$ and $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$.
0:Plaintext sequence: $P=(P_{1},P_{2},...,P_{K})$.
[…]
6:for $i=K;i\geq 1;i$-- do
7:  $(x_{i},S)\leftarrow\mathbb{D}_{w_{0},w_{1}}(C_{2i},S)$
8:  $S\leftarrow(S\oplus(1\&x_{i}))$
9:  $(P_{i},S)\leftarrow\mathbb{D}_{w_{0},w_{1}}(C_{2i-1},S)$
10:  $S\leftarrow(S\oplus(1\&P_{i}))$
11:end for
4.2.4 Basic analysis of EagleEagle
It should also be noted that after being encrypted by the $Eagle$ encryption algorithm, the length of the ciphertext is about 2 times the length of the plaintext. This is because in the preparation stage of the $Eagle$ encryption algorithm, independent random numbers with a length of about 1 times the plaintext are generated, and then these random numbers are encrypted using the known short key. The encrypted ciphertext formed by these random numbers is also bound to the final ciphertexts.
4.2.5 A construction of one-way function
When given the ciphertext ($C_{1}$, $C_{2}$, …, $C_{2K}$) and $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$ of the $Eagle$ encryption algorithm, we define the following function.
$f(w_{0},w_{1},w_{2},w_{3})=(P_{1},P_{2},...,P_{K})$ (4.1)
where $(w_{0},w_{1},w_{2},w_{3})$ is the short key and $(P_{1},P_{2},...,P_{K})$ is the plaintext decrypted by the $Eagle$ decryption algorithm.
The inverse function of $f$ can be defined as
$f^{-1}(P_{1},P_{2},.,P_{K})=\{(w_{0},w_{1},w_{2},w_{3})|$ (4.2)
$f(w_{0},w_{1},w_{2},w_{3})=(P_{1},P_{2},.,P_{K})\}$
Solving $f^{-1}$ is equivalent to finding the key $(w_{0},w_{1},w_{2},w_{3})$ to match the plaintext $(P_{1},P_{2},...,P_{K})$.
[…]
Intuitively, the Eagle encryption algorithm is composed of two independent parts, each consisting of its own key, plaintext, and ciphertext. Given the correspondence between the plaintext ($P_{1},P_{2},...,P_{K}$) and ciphertext ($C_{1}$, $C_{2}$, …, $C_{2K}$), $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$ of the Eagle encryption algorithm, the problem of solving the key $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3})$ is equivalent to the problem of solving the respective keys in two independent encryption algorithms under the condition that only their plaintext or ciphertext is known. The first encryption algorithm solves its key $(w_{0},w_{1})$ under the condition that its plaintext $(P_{1},P_{2},...,P_{K})$ and its ciphertext $(C_{1},C_{2},...,C_{2K})$ are known. And in another encryption algorithm, the key $(w_{2},w_{3})$ need to be solved under the condition that only the ciphertext ($C_{1}^{{}^{\prime}}$, $C_{2}^{{}^{\prime}}$, $S_{2}^{{}^{\prime}}$) is known.
In fact, for these two independent encryption algorithms, under the condition of only knowing their respective plaintext or ciphertext, every possible key cannot be reasonably excluded, and every possible key must be exhaustively searched. From the perspective of the entire Eagle encryption algorithm, at least one set of keys from $(w_{0}$, $w_{1})$ or $(w_{2}$, $w_{3})$ needs to be exhaustively searched to find a valid key. The computational complexity of exhausting either $(w_{0}$, $w_{1})$ or $(w_{2}$, $w_{3})$ is exponentially.
Without specifying any intermediate state $S$ or $y_{0}$, the problem of solving $f^{-1}$ is equivalently transformed into solving a system of equations about the unknown variables as $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3}$) and $(x_{1},x_{2},...,x_{K})$ and $y_{1}$, Any equation in this system of equations can be regarded as an identity equation about the unknown variables, and it is impossible to find an effective solution.
On the other hand, given $(P_{1}$, $P_{2}$, … , $P_{K})$ and $(C_{1}$, $C_{2}$, …, $C_{2K})$ and $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$, every intermediate state $S$ can take any value, the computational complexity of traversing every intermediate state is $O(2^{L})$.
This indicates that the computational complexity of solving $f^{-1}$ is at least $O(2^{L})$.
Let’s give the following theorem.
[Theorem 5] When given the ciphertext ($C_{1}$, $C_{2}$, …, $C_{2K})$, $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}}$) and the plaintext $(P_{1},P_{2},...,P_{K})$ of the $Eagle$ encryption algorithm, the computational complexity of the problem of cracking the key $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3})$ is at least $O(2^{L})$.
Proof.
The entire proof process is divided into the following two parts.
(i) Any intermediate state $S$ is unknown
By observing the structure of Algorithm 3 and Algorithm 4, without knowing any $S$, the problem of cracking the key $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3})$ is equivalent to solving the following system of equations
[…]
Where * are placeholder. $(P_{1}$, $P_{2}$, …, $P_{K})$, $(C_{1}$, $C_{2}$, …, $C_{2K})$, $(C_{1}^{{}^{\prime}}$, $C_{2}^{{}^{\prime}}$, $S_{2}^{{}^{\prime}})$ are known variables. $(w_{0},w_{1},w_{2},w_{3})$, $(x_{1},x_{2},...,x_{K})$, $y_{1}$ are unknown variables to be solved.
According to the Theorem 3, any variable in any equation of the system of equations (4.3) and (4.4) can take any value, so each equation can be regarded as an identity equation.
For the system of equations (4.5), the first equation is also an identity. The unknown variable $y_{1}$ in the second equation is a free variable, this equation holds for any $(w_{2},w_{3})$, thus we need to traverse all $(w_{2},w_{3})$ to solve this equation.
We can conclude that the computational complexity of the problem of cracking the key $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3})$ is at least $O(2^{2L-1})$.
(ii) One of the intermediate states $S$ is known.
Without loss of generality we suppose $S_{1}$ is known,according to the Theorem 2, $S_{1}$ depends only on $P_{1}$ and $(w_{0},w_{1})$. For any given combination of $S_{1}$ and $P_{1}$, we can always find some reasonable $(w_{0},w_{1})$, which requires a conclusion that we haven’t yet demonstrated, we will provide a detailed argument in the next chapter, see Theorem 7 for details.
Since we can find some reasonable $(w_{0},w_{1})$ for any given $S_{1}$, thus we need to traverse all $S_{1}$ to solve $(w_{0},w_{1})$ and $(w_{2},w_{3})$.
We can conclude that the computational complexity of the problem of cracking the key $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3})$ is at least $O(2^{L})$.
[Theorem 6] The function $f$ defined in (4.1) is a one-way function.
[…]
5 How to achieve only exhaustive search to crack keys
5.1 In depth analysis of Eagle encoding algorithm
@@
[Theroem 7] $\forall m,s$, there exists $2^{L-1}$ different groups $(w_{0},w_{1})$ satisfy that $\mathbb{E}_{w_{0},w_{1}}(m,*)=(*,s)$. For any $\delta$, there exists a unique $w_{0}$ that $\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m,*)=(*,s)$.
@@
From (5.2), we know $\mathbb{E}_{w_{0},w_{1}}(m,*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m,\delta))$, through a simple transformation we obtain $w_{0}^{+(L-1)}=g_{L}(m,\delta)\oplus s$, it is obvious that for any $\delta,m,s$, the right side of the equation is a difinite result, so there exists a unique $w_{0}$ valid.
@@
The conclusion of Theorem 7 indicates that, for any $\mathbb{E}_{w_{0},w_{1}}(m,*)=(*,s)$, when only given $(m,s)$, there are exponentially different $(w_{0},w_{1})$ valid.
[Theorem 8] For any $m_{1},s_{1},m_{2},s_{2}$, when the parity between $m_{1}$ and $m_{2}$ is different, there exists a unique $(\delta,w_{0})$ that $\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1},*)=(*,s_{1})$ and $\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2},*)=(*,s_{2})$.
@@
The conclusion of Theorem 8 indicates that, for any $\mathbb{E}_{w_{0},w_{1}}(m_{1},*)=(*,s_{1})$ and $\mathbb{E}_{w_{0},w_{1}}(m_{2},*)=(*,s_{2})$, when only given $m_{1},s_{1},m_{2},s_{2}$, there is a 50% probability that a unique $w_{0},w_{1}$ valid.
@@
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{1},\delta))$ (5.3)
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{2},\delta))$
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1}^{{}^{\prime}},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{1}^{{}^{\prime}},\delta))$
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2}^{{}^{\prime}},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{2}^{{}^{\prime}},\delta))$
@@
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m\oplus\Delta m,*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m,\delta)\oplus j(\delta,\Delta m))$
@@
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1}^{{}^{\prime}},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{1},\delta)\oplus j(\delta,\Delta m))$ (5.4)
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2}^{{}^{\prime}},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{2},\delta)\oplus j(\delta,\Delta m))$
@@
[Theorem 10] For the function $g_{L}(m,\delta)$ in (5.2), fix any $\delta$, where $\delta$ has an odd number of bits that are 1. Fix any $u$, there exists a unique $m$ that $g_{L}(m,\delta)=u$.
@@
Fix any $\delta$, set $w_{0}\in\{0,1\}^{L},w_{1}=w_{0}\oplus\delta$, excute Eagle encoding algorithm for $m$, we obtain $\mathbb{E}_{w_{0},w_{1}}(m,*)=(*,s)$, then $u=s\oplus w_{0}^{+L}$, $g_{L}(m,\delta)=u$, thus $m\to u$ is monomorphism.
On the other hand, fix any $\delta$, set $w_{0}\in\{0,1\}^{L},w_{1}=w_{0}\oplus\delta$, excute Eagle decoding algorithm for $s=u\oplus w_{0}^{+L}$, we obtain $\mathbb{D}_{w_{0},w_{1}}(*,s)=(m,*)$, thus $u\to m$ is monomorphism.

### 06 Apr 2026 v8

@@ 1 Introduction
In Chapter 3, we propose a novel encoding mechanism which we named Eagle. The distinguishing feature of Eagle encoding algorithm from any other encoding algorithm is that for any given group of input and output texts, the key parameters are unsolvable.
In Chapter 4, we further delve into the mathematical laws behind the Eagle encoding algorithm and discover an operation that satisfies the allocation law for XOR operations. This operation is very similar to the allocation law of multiplication operations, but does not satisfy the commutative and associative laws, we refer to it as a “right multiplication operation on binary”. Then we construct a binary operation of two-dimensional vectors which serves as a key ingredient for the subsequent design of encryption algorithms.
@@
In fact the encryption algorithm developed in Chapter 5 does not meet the security requirements for the key. This is because, when knowing the plaintext-ciphertext correspondence, although the complexity of cracking the key is exponential to the length of the key, it does not guarantee that only exhaustive search is effective for cracking the key. There may be some other attack methods that can effectively reduce the difficulty of cracking the key. Such as for a 256 bit key, although it can ensure that the computational complexity of cracking the key is at least $O(2^{64})$, this is still far from the ideal security of $O(2^{256})$. The encryption algorithm constructed in Chapter 5 only analyzes the computational complexity of cracking the key given a plaintext ciphertext correspondence, which does not guarantee its ability to resist linear and differential attacks, For attackers, they can construct plaintexts with certain characteristics multiple times, then find some structures that we did not anticipate from the relationship between plaintext and ciphertext.
In Chapter 6, we first construct nonlinear operations of one parameter on another, such as “high-order” operations. Based on this nonlinear construction, we redesigned a block symmetric encryption algorithm and named it $Eagle^{*}$. For $Eagle^{*}$, knowing any plaintext ciphertext correspondence, there is no more effective way for attackers to crack the key than exhaustive search.
@@ 2 Notations
$\mathbb{E}_{w_{0},w_{1}}(m)$ - Excute Eagle encoding algorithm as shown in Algorithm 1.
$\mathbb{D}_{w_{0},w_{1}}(s)$ - Excute Eagle decoding algorithm as shown in Algorithm 2.
$m*\delta$ - Excute the right multiplication of $m$ with $\delta$ as shown in Algorithm 3.
$s/\delta$ - Excute the right division of $s$ to $\delta$ as shown in Algorithm 4.
$m*\delta^{2}$ - Excute the right multiplication of $m$ with $\delta$ twice as $m*\delta*\delta$ as shown in Algorithm 7.
$s/\delta^{2}$ - Excute the right division of $s$ to $\delta$ twice as $s/\delta/\delta$ as shown in Algorithm 8.
$(m_{1},m_{2})\bigotimes(k_{1},k_{2})=(c_{1},c_{2})$ - A binary operation of two-dimensional vectors defined in chapter 4.2.
@@ 3 Introduction to Eagle encoding algorithm
$\mathbb{E}_{w_{0},w_{1}}(m)=s_{L}$ : use the key $(w_{0},w_{1})$ to execute Eagle encoding algorithm (Algorithm 1) on any initial state $s_{0}$ and the given input $m$ to obtain $c$ and $s_{L}$, which $s_{0}$ and $c$ will not be analyzed anymore.
$\mathbb{D}_{w_{0},w_{1}}(s_{L})=m$ : use the key $(w_{0},w_{1})$ to execute Eagle decoding algorithm (Algorithm 2) on any $c$ and the given $s_{L}$ to obtain $s_{0}$ and $m$.
@@
4 In depth analysis of Eagle encoding algorithm
@@
[Theroem 3] $\forall m,s$, there exists $2^{L-1}$ different groups $(w_{0},w_{1})$ satisfy that $\mathbb{E}_{w_{0},w_{1}}(m)=s$. For any $\delta$, there exists a unique $w_{0}$ that $\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m)=s$.
@@
From (4.2), we know $\mathbb{E}_{w_{0},w_{1}}(m)=w_{0}^{+(L-1)}\oplus g_{L}(m,\delta)$, through a simple transformation we obtain $w_{0}^{+(L-1)}=g_{L}(m,\delta)\oplus s$, it is obvious that for any $\delta,m,s$, the right side of the equation is a difinite result, so there exists a unique $w_{0}$ valid.
@@
The conclusion of Theorem 3 indicates that, for any $\mathbb{E}_{w_{0},w_{1}}(m)=s$, when only given $(m,s)$, there are exponentially different $(w_{0},w_{1})$ valid.
[Theorem 4] For any $m_{1},s_{1},m_{2},s_{2}$, when the parity between $m_{1}$ and $m_{2}$ is different, there exists a unique $(\delta,w_{0})$ that $\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1})=s_{1}$ and $\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2})=s_{2}$.
@@
The conclusion of Theorem 4 indicates that, for any $\mathbb{E}_{w_{0},w_{1}}(m_{1})=s_{1}$ and $\mathbb{E}_{w_{0},w_{1}}(m_{2})=s_{2}$, when only given $m_{1},s_{1},m_{2},s_{2}$, there is a 50% probability that a unique $w_{0},w_{1}$ valid.
@@
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1})=w_{0}^{+(L-1)}\oplus g_{L}(m_{1},\delta)$ (4.3)
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2})=w_{0}^{+(L-1)}\oplus g_{L}(m_{2},\delta)$
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1}^{{}^{\prime}})=w_{0}^{+(L-1)}\oplus g_{L}(m_{1}^{{}^{\prime}},\delta)$
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2}^{{}^{\prime}})=w_{0}^{+(L-1)}\oplus g_{L}(m_{2}^{{}^{\prime}},\delta)$
@@
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m\oplus\Delta m)=w_{0}^{+(L-1)}\oplus g_{L}(m,\delta)\oplus j(\delta,\Delta m)$
@@
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1}^{{}^{\prime}})=w_{0}^{+(L-1)}\oplus g_{L}(m_{1},\delta)\oplus j(\delta,\Delta m)$ (4.4)
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2}^{{}^{\prime}})=w_{0}^{+(L-1)}\oplus g_{L}(m_{2},\delta)\oplus j(\delta,\Delta m)$
@@
[Theorem 6] For the function $g_{L}(m,\delta)$ in (4.2), fix any $\delta$, where $\delta$ contains an odd number of bits that are 1. Fix any $u$, there exists a unique $m$ that $g_{L}(m,\delta)=u$.
@@
Fix any $\delta$, set $w_{0}\in\{0,1\}^{L},w_{1}=w_{0}\oplus\delta$, excute Eagle encoding algorithm for $m$, we obtain $\mathbb{E}_{w_{0},w_{1}}(m)=s$, then $u=s\oplus w_{0}^{+L}$, $g_{L}(m,\delta)=u$, thus $m\to u$ is monomorphism.
On the other hand, fix any $\delta$, set $w_{0}\in\{0,1\}^{L},w_{1}=w_{0}\oplus\delta$, excute Eagle decoding algorithm for $s=u\oplus w_{0}^{+L}$, we obtain $\mathbb{D}_{w_{0},w_{1}}(s)=m$, thus $u\to m$ is monomorphism.